= O GEOMETRY Surface area of a cylinder Find the surface area of a cylinder with a base diameter of 8 ft and a height of 5 ftHelp please
Answer #1

Surface area of a cylinder equals to π.d.h

Where,

d is diameter of base , and h is height of cylinder.

Therefore, Surface area = π × 8 × 5 = 40 × 3.14 = 125.6

Surface area equals to 125.6 square feet.